CFSE’1 Rennes, 8 juin - 11 juin 1999

1ère Conférence Française sur les Systèmes d’Exploitation

Authors Guidelines Recommandations aux auteurs

Kermarrec, A.-M. IRISA, Campus de Beaulieu F-35042 Rennes Cedex

Résumé

In this document, we present authors guidelines for submissions to CFSE’1. Submitted papers should respect this format to be integrated in the conference proceedings.Titles must be available in french and in english.

Abstract

Abstracts in french and english are compulsory.

1 Guidelines

This document has been done with LaTex2e and the latex style Cfse1.sty. If you can not use LaTex2e, we ask you to respect a similar format. Papers will be published in this format (no reduction). Pages must not be numbered.

Sections must be numbered. We ask you not to change the page size or the fonts used (11pt for the text). Papers must be no longer than 12 pages including bibliography, graphics and appendix. Official languages of the conference are French and English. French-speaking authors must submit in French.

2 Submissions

Electronic submissions are encouraged (postscript or PDF formats only) and must be sent at the e-mail address cfse1@irisa.fr. The title of the paper, the contact author, his/her acomplete address and e-mail, the abstract of the paper and a list of key words must be sent in ASCII with the submission.

Authors who don’t have access to an e-mail may send their submission by snail mail at the following address:

Valérie Issarny
Soumission à CFSE'1
IRISA, Campus de Beaulieu
F-35042 Rennes Cedex

Outstanding papers will be eligible for publication in a special issue of “Calcul Parallele, Reseaux et Systemes”.

3 Rappel des principales dates

– Deadline for receipt of submissions: January15, 1999

– Acceptance notification: March 15, 1999

– Final programme and call for participation: April 1st, 1999 – Deadline for camera-ready papers: April 15, 1999

– Cfse’1 in Rennes: June 8-11, 1999